# BigTokDetect: A Clinically-Informed Vision–Language Modeling Framework for Detecting Pro-Bigorexia Videos on TikTok

### v2

@@ Abstract
Social media platforms increasingly struggle to detect harmful content that promotes muscle dysmorphic behaviors, particularly pro-bigorexia content that disproportionately affects adolescent males. Unlike traditional eating disorder detection focused on the "thin ideal," pro-bigorexia material masquerades as legitimate fitness content through complex multimodal combinations of visual displays, coded language, and motivational messaging that evade text-based detection systems. We address this challenge by developing BigTokDetect, a clinically-informed detection framework for identifying pro-bigorexia content on TikTok. We introduce BigTok, the first expert-annotated multimodal dataset of over 2,200 TikTok videos labeled by clinical psychologists and psychiatrists across five primary categories spanning body image, nutrition, exercise, supplements, and masculinity. Through comprehensive evaluation of state-of-the-art vision language models, we achieve 82.9% accuracy on primary category classification and 69.0% on subcategory detection via domain-specific finetuning. Our ablation studies demonstrate that multimodal fusion improves performance by 5-10% over text-only approaches, with video features providing the most discriminative signals. These findings establish new benchmarks for multimodal harmful content detection and provide both the computational tools and methodological framework needed for scalable content moderation in specialized mental health domains.
BigTokDetect: A Clinically-Informed Vision–Language Modeling Framework for Detecting Pro-Bigorexia Videos on TikTok
@@ 1 Introduction
Social media platforms face mounting challenges in detecting harmful content that significantly impacts user mental health Chancellor and Choudhury (2020); Gorwa et al. (2020). Contemporary platforms like TikTok algorithmically amplify content promoting unrealistic body ideals Becker (2004); Minadeo and Pope (2022), yet much harmful material exists in gray areas where legitimate discussions intersect with dangerous messaging Gillespie (2018). These narratives emerge through complex multimodal combinations—visual imagery, audio cues, and textual descriptions—that traditional text-based detection systems cannot adequately capture Kiela et al. (2020a, b). Effective automated moderation requires sophisticated multimodal approaches that can parse nuanced signals across video, audio, and text Gimeno-Gómez et al. (2024).
Pro-bigorexia content—material promoting muscle dysmorphic behaviors—exemplifies these multimodal detection challenges at their most complex. This harmful content promotes compulsive muscle-building behaviors that disproportionately affect adolescent males Pope Jr et al. (1997); Mitchison et al. (2022), yet remains severely understudied in computational research compared to traditional eating disorder detection Chancellor and Choudhury (2020). Pro-bigorexia material masquerades as legitimate fitness content through subtle combinations of muscular displays, extreme workout demonstrations, and coded supplement language Murray et al. (2017); Kamkari (2025). The visual similarity between harmful and beneficial fitness content, combined with evolving coded terminology, renders current text-based detection approaches fundamentally inadequate.
Current automated detection systems face critical limitations when confronting such content. Foundation models trained on general corpora lack domain-specific knowledge to recognize subtle clinical markers and euphemistic language patterns Murray et al. (2017). Existing systems rely heavily on keyword filtering and text-based signals, missing critical visual and behavioral cues embedded in video content Kiela et al. (2020b); Gorwa et al. (2020). The dynamic nature of social media—where creators deliberately evolve language to evade detection—further challenges automated approaches Gillespie (2018). Effective detection requires integrated analysis of visual behavioral cues, audio transcripts, and textual context—capabilities that remain underexplored in computational mental health research Gimeno-Gómez et al. (2024).
@@
  • We develop the first clinically-informed taxonomy for automated pro-bigorexia detection, establishing fine-grained categories spanning body image, nutrition, supplement abuse, exercise practices, and masculinity that enable systematic computational analysis.
  • We introduce BigTok, a multimodal dataset of over 2,200 expert-annotated TikTok videos that enables robust evaluation of vision-language models on challenging harmful content detection tasks.
  • We construct BigTokDetect, a detection framework that achieves state-of-the-art (SOTA) performance using proprietary and open-source VLMs, with our best models reaching 95.1% accuracy on primary category classification and 91.0% on fine-grained subcategory detection through fine-tuning.
  • We demonstrate through comprehensive ablation studies that multimodal fusion improves detection performance by 5-10% over text-only approaches, with video features providing the most discriminative signals for identifying subtle pro-bigorexia behaviors.
Figure 1 summarizes our comprehensive approach to multimodal pro-bigorexia detection. Our methodology demonstrates that clinically-informed vision-language models can effectively identify subtle harmful content that evades traditional detection systems. Through systematic evaluation across multiple model architectures and training paradigms, we achieve substantial improvements over existing text-based methods, establishing a new paradigm for detecting nuanced harmful content on social media platforms. The BigTokDetect framework provides both the computational tools and replicable methodology needed to address the growing challenge of automated content moderation in mental health domains.
@@ 2 Related Work
Muscle dysmorphic disorder, or "bigorexia," involves preoccupation with insufficient muscularity, driving compulsive behaviors including exercise, rigid dieting, and supplement abuse Pope Jr et al. (1997); Cooper et al. (2020). Frequent exposure to muscularity-oriented TikTok and Instagram content—including body-transformation videos, supplement promotions, and steroid use—has been linked to elevated rates of probable muscle dysmorphia Mitchison et al. (2022); Ganson et al. (2025). Hypermasculine online subcultures (e.g., the "manosphere") intensify these pressures through toxic social comparisons and body-optimization trends Kamkari (2025). Despite this prevalence, both psychological research and platform moderation remain focused on anorexia and bulimia, leaving bigorexia largely unmoderated Lookingbill et al. (2023).
@@ 3 Pro-bigorexia Content Taxonomy
We develop an empirically grounded taxonomy for pro-bigorexia content characterization, following an iterative four-step process. First, we compile a seed list of keywords from fitness, bodybuilding, exercise, and diet domains, supplemented by terms from foundational muscle dysmorphia research and popular media reports Cafri et al. (2008); Sánchez et al. (2024); Willett (2023). Then, using these keywords, we sample 500 TikTok videos over 40 hours, and open-code emergent patterns of bigorexia, such as body-check demonstrations, supplement endorsements, extreme workout narratives, diet tutorials). Next, we draft preliminary categories based on the DSM-5 clinical criteria American Psychiatric Association (2013), validated measures such as the Drive for Muscularity Scale McCreary and Sasse (2000) and the Muscle Dysmorphic Disorder Inventory Zeeck et al. (2018). Finally, we convene with clinical experts specializing in eating disorders and body image psychopathologies to refine categories, clarify specialized or coded language (especially around steroids and supplements), and disambiguate edge cases. We iterate to discover new potential categories and keywords.
Our final taxonomy for pro-bigorexia TikTok content consists of five primary categories: Relationship to Body, Relationship to Food, Relationship to Exercise, Supplements, and Masculinity. Each is further divided into clinically grounded subcategories (Table 12, Appendix A.1). For each category and subcategory, we provide clear definitions, representative keywords, practical guidelines for recognizing problematic content, and sample videos. This approach yields a consistent, multimodal annotation of overt and implicit signals of pro-bigorexia behaviors.
@@ 4.1 Data Collection
We source our corpus from TikTok via its official API TikTok (2025). We query videos from January 2019 to January 2025 to capture pre- and post-pandemic trends. Guided by domain experts, we curate 40 high-precision query terms mapped to taxonomy subcategories (Table 12, Appendix A.1), retrieving up to 1,000 videos and their metadata per term. We randomly sample videos per group of keywords that belong to each primary and subcategory to 2,400 videos for annotation. Only video content and captions are exposed to annotators to protect user privacy (anonymization steps are described in Appendix C). We also compile negative (irrelevant) examples using 42 trending hashtags from TikTok’s Creative Center TikTok Creative Center , following established supervised classification practices Kiela et al. (2020b).
@@ 4.2 Annotating Bigorexia Content
We base our annotation interface on Amazon Mechanical Turk via an invitation-only pool (Appendix B.3). Annotators are instructed to select the first (mandatory) primary–subcategory pair ($t_{1},s_{1}]$) and the second (optional) primary–subcategory pair ($t_{2},s_{2}]$). Every video is also rated for harm on a discrete 5-point Likert scale with 0.5-point increments (1=not harmful and 5=very harmful). The annotation interface is shown in Appendix B.2.
@@
As a pilot study, each annotator independently labels a set of 20-30 videos to validate taxonomy coverage, identify potential edge cases, and familiarize themselves with the annotation interface. We monitor disagreements and encourage detailed note-taking. Annotators then join a group session to discuss flagged issues, collaboratively annotate selected videos while explaining their reasoning, until consensus is reached.
@@
We then flag examples with weak agreement or disagreement on pro-bigorexia detection, and examples where severity scores differ by more than two points. The two annotators are invited to a structured discussion to articulate their reasoning, review cases that need consultation, address disagreements, and eventually aim to reach consensus, although consensus is not mandatory. After the discussion, each annotator independently re-annotates the flagged videos. Further details on this procedure are provided in Appendix  B.4.
@@
While each video may receive multiple label pairs (primary-subcategory), for sake of easy evaluation, we assign a single label pair to each video. For videos with perfect agreement, we randomly choose of the label pairs. For strong agreement, we adopt the shared primary–subcategory pair; for weak agreement, we choose the mandatory (first) label over the optional (second) one; remaining disagreements are resolved by random tie-breaking. The final harm severity score is the average of the two annotators’ ratings. This simplification reflects current limitations of LLMs in reliable multi-label classification Ma et al. (2025).
We assess inter-rater reliability with Cohen’s $\kappa$  Cohen (1960) for two annotators. Initial agreement is moderate (strong $\kappa=0.43$$0.69$; weak $\kappa=0.59$$0.83$; Table 6). After structured consensus discussions and re-annotation, reliability increases substantially (strong $\kappa=0.58$$0.81$; weak $\kappa=0.78$$0.94$), ensuring that our iterative process yields high-quality, consistent labels. All Cohen’s $\kappa$ values for each batch are shown in Table 6 (Appendix B.5)
@@ 4.3 Data Statistics
Primary Category Count
Relationship to Body 449
Relationship to Exercise 450
Relationship to Food 317
Supplement Abuse 220
Relationship to Masculinity 131
Irrelevant 639
Subcategory Count
Muscularity Self-objectification 278
Leanness Self-objectification 72
Muscle Dissatisfaction 78
Rigid Food Rules 136
Unsafe Food 79
Cheat Meals 74
Excessive Exercise 109
Predebting Exercise 22
Maladaptive Coping 89
Exercise-Induced Functional Impairment 34
Toxic Motivation 107
Anabolic Steroids 68
Legal APEDs 50
Hormone Therapy 82
Relationship to Masculinity 131
Other 103
Irrelevant 639
Table 1: Distribution of categories and subcategories BigTok. We include the primary categories Irrelevant and Relationship to Masculinity, which can also be considered as subcategories.
Our final BigTok dataset comprises 2,210 annotated TikTok videos. The median video duration is 20.6 seconds, and 59.8% of videos are under 30 seconds. Table 1 shows the label distribution. At the primary category level, excluding Irrelevant, Relationship to Exercise (450) is the most dominant class. At the subcategory level, Muscularity Self-objectification predominates (278), reflecting the popularity of the genre of videos in which people display their muscles. The Other subcategory covers videos within a primary category that do not match predefined subcategories (e.g., humor or lifestyle clips); these are excluded from subcategory-level benchmarking. Some examples from BigTok are shown in Tables 13, 14, and 15 Appendix A.1. Additional annotation statistics are provided in Appendix B.5.
@@ 5 Classifying Multimodal Pro-Bigorexia Content via VLMs
To develop an automated detection system for pro-bigorexia content, we apply SOTA VLMs to this challenging multimodal classification task. We leverage both commercial and open-source VLMs through zero-shot prompting, few-shot learning, and fine-tuning approaches to achieve high performance in identifying subtle pro-bigorexia behaviors across video, audio, and text modalities.
@@ 5.2 Models
Given the multimodal nature of our classification task, involving text from captions, audio transcripts, images, and video content, we require SOTA models designed for multimodal data, particularly video understanding. We evaluate three commercial API-based VLMs: GPT-4.1 OpenAI (2025) and Claude-Sonnet-4 Anthropic PBC (2025) (both process images), while Gemini-2.5-Flash Comanici et al. (2025) natively accepts video input. We also evaluate two open-source models: Qwen2.5-VL Wang et al. (2024), and InternVL3 Zhu et al. (2025), which uses Qwen2.5 pre-trained base models and Variable Visual Position, which achieves strong performance on video benchmarks.
Model sizes and versions are listed in Table 10, Appendix D.1. Due to computational constraints, we focus on small and medium-sized models. For open-source models, we set temperature = 0.1 and top_p = 0.9. For API models, we retain the default hyperparameters temperature = 1.0 and top_p = 1.0. Implementation details and hardware specifications are provided in Appendix D.1. Details about the train and test datasets are in Section C, Appendix.

### v3

# BigTokDetect: A Clinically‑Informed Vision–Language Modeling Framework for Detecting Pro‑Bigorexia Videos on TikTok

@@ Abstract
Social media platforms face escalating challenges in detecting harmful content promoting muscle dysmorphic behaviors and cognitions (bigorexia), which disproportionately affects adolescent males and evades moderation by camouflaging as legitimate fitness advice. We address this challenge with BigTokDetect, a clinically informed framework for identifying pro-bigorexia content on TikTok. We introduce BigTok, the first expert-annotated multimodal benchmark dataset of over 2,200 TikTok videos labeled by clinical psychiatrists across five categories and eighteen fine-grained subcategories. Comprehensive evaluation of state-of-the-art vision-language models (VLMs) reveals that while commercial zero-shot models achieve the highest accuracy on broad primary categories, supervised finetuning allows smaller open-source models to achieve better fine-grained subcategory detection. Ablation studies demonstrate that multimodal fusion improves performance by 5-15%, with video features providing the most discriminative signals. These findings support a grounded moderation approach, automating detection of explicit harms while flagging ambiguous content for human review, and establish a scalable framework for harm mitigation in emerging mental health domains.
BigTokDetect: A Clinically‑Informed Vision–Language Modeling Framework for Detecting Pro‑Bigorexia Videos on TikTok
@@ 1 Introduction
Social media platforms face mounting challenges in detecting harmful content that significantly impacts user mental health Chancellor and Choudhury (2020); Gorwa et al. (2020). Contemporary platforms like TikTok algorithmically amplify content promoting unrealistic body ideals Becker (2004); Minadeo and Pope (2022), yet much of the harmful material exists in gray areas where legitimate discussions intersect with unsafe messages Gillespie (2018). These narratives emerge through complex multimodal combinations, which include visual imagery, audio cues, and textual descriptions that traditional text-based detection systems do not adequately capture Kiela et al. (2020a, b). Effective automated moderation requires comprehensive multimodal approaches that can parse nuanced signals across video, audio, and text gimenogómez2024readingframesmultimodaldepression.
Pro-bigorexia content, material promoting muscularity as an aspirational goal, exemplifies these multimodal detection challenges. This content promotes compulsive muscle-building behaviors that disproportionately affect adolescent males Pope et al. (1997); Mitchison et al. (2022), yet remains severely understudied in computational research compared to traditional eating disorder detection Chancellor and Choudhury (2020). Pro-bigorexia material often masquerades as legitimate fitness content through subtle combinations of muscularity displays, extreme workout demonstrations, and coded supplement language Murray et al. (2017); Kamkari (2025). The visual similarity between harmful and beneficial fitness content, combined with evolving jargon, renders current detection approaches fundamentally inadequate.
Current automated detection systems face critical limitations while moderating such content. Foundation models trained on general corpora lack domain-specific knowledge to recognize subtle clinical markers and coded language patterns Murray et al. (2017). Systems relying on keyword filtering and text-based signals miss critical visual and behavioral cues embedded in video content Kiela et al. (2020b); Gorwa et al. (2020). The dynamic nature of social media, where creators deliberately evolve language to evade detection, poses another challenge to automated approaches Gillespie (2018). Effective detection requires integrated analysis of visual behavioral cues, audio transcripts, and textual context, capabilities that remain underexplored in computational mental health research gimenogómez2024readingframesmultimodaldepression.
@@
  • We develop the first clinically-informed taxonomy for automated detection of pro-bigorexia content, establishing fine-grained categories spanning body image, nutrition, supplement abuse, exercise practices, and masculinity that enable systematic computational analysis.
  • We introduce BigTok, a multimodal benchmark dataset of over 2,200 expert-annotated TikTok videos that enables robust evaluation of vision-language models (VLMs) on challenging harmful content detection tasks.
  • We construct BigTokDetect, a detection framework that achieves state-of-the-art (SOTA) performance using proprietary and open-source VLMs, with our best models reaching 82.9% accuracy on primary category classification and 69.0% on fine-grained subcategory detection through finetuning.
  • We demonstrate through comprehensive ablation studies that multimodal fusion improves detection performance by 5-15% over text-only approaches, with video features providing the most discriminative signals for identifying subtle pro-bigorexia cues.
Figure 1 summarizes our comprehensive approach to multimodal pro-bigorexia detection. Our results demonstrate that clinically-informed VLMs can effectively identify subtle harmful content that evades traditional detection systems. Through systematic evaluation across multiple model architectures and training paradigms, we achieve substantial improvements over existing methods, establishing a new paradigm for detecting nuanced harmful content on social media platforms. The BigTokDetect framework provides both the computational tools and reproducible methodology needed to address the growing challenge of automated content moderation in mental health domains.
@@ 2 Related Work
Muscle dysmorphic disorder, or “bigorexia,” involves preoccupation with insufficient muscularity, driving compulsive behaviors including exercise, rigid dieting, and supplement abuse Pope et al. (1997); Cooper et al. (2020). Frequent exposure to muscularity-oriented TikTok and Instagram content, including body-transformation videos, supplement promotions, and steroid use, has been linked to elevated rates of probable muscle dysmorphia Mitchison et al. (2022); Ganson et al. (2025). Hypermasculine online subcultures (e.g., the “manosphere”) intensify these pressures through toxic social comparisons and body-optimization trends Kamkari (2025). Despite this prevalence, both psychological research and platform moderation remain focused on anorexia and bulimia, leaving bigorexia largely unmoderated Lookingbill et al. (2023).
@@ 3 Pro-bigorexia Content Taxonomy
We develop an empirically grounded taxonomy for pro-bigorexia content characterization, following an iterative four-step process. First, we compile a seed list of keywords from fitness, bodybuilding, exercise, and diet domains, supplemented by terms from foundational muscle dysmorphia research and popular media reports Cafri et al. (2008); Chu et al. (2024); Willett (2023). Then, using these keywords, we retrieve 500 TikTok videos over 40 hours, and open-code emergent patterns of bigorexia, such as body-check demonstrations, supplement endorsements, extreme workout narratives, diet tutorials). Next, we draft preliminary categories based on the DSM-5 clinical criteria American Psychiatric Association (2013), validated measures such as the Drive for Muscularity Scale McCreary and Sasse (2000) and the Muscle Dysmorphic Disorder Inventory Zeeck et al. (2018). Finally, we convene with clinical experts specializing in eating disorders and body image psychopathologies to refine categories, clarify specialized or coded language (especially around steroids and supplements), and disambiguate edge cases. We iterate to discover new potential categories and keywords.
Our final taxonomy for pro-bigorexia TikTok content consists of five primary categories: Relationship to Body, Relationship to Food, Relationship to Exercise, Supplements, and Masculinity (Table 7, Appendix A.1). Each is further divided into clinically grounded subcategories. For each category and subcategory, we provide clear definitions, representative keywords, practical guidelines for recognizing problematic content, and sample videos. The taxonomy facilitates a consistent, multimodal annotation of overt and implicit signals of pro-bigorexia behaviors.
@@ 4.1 Data Collection
We source our corpus from TikTok via its official API TikTok (2025), querying videos from January 2019 to January 2025 to capture pre- and post-pandemic trends. Guided by domain experts, we curate 40 high-precision query terms mapped to taxonomy subcategories (Table 7, Appendix A.1), retrieving up to 1,000 videos per term, from which we randomly sample 2,400 videos for annotation. To protect user privacy, only video content and captions are exposed to annotators (anonymization steps are described in Appendix C). To ensure robust supervised classification, we construct a negative control group (the Irrelevant category) via two mechanisms: (1) proactive sampling using 42 generic trending hashtags (e.g., #viral, #travel) from TikTok’s Creative Center TikTok Creative Center ; and (2) re-assigning videos flagged by experts during annotation as unrelated to bigorexia. This strategy ensures the model learns to distinguish harmful content from both general entertainment and benign daily activities, following established supervised classification practices Kiela et al. (2020b).
@@ 4.2 Annotating Bigorexia Content
We base our annotation interface on Amazon Mechanical Turk via an invitation‐only pool (Appendix B.3). Annotators are instructed to select the first (mandatory) primary–subcategory pair ($t_{1},s_{1}]$) and the second (optional) primary–subcategory pair ($t_{2},s_{2}]$). Every video is also rated for harm on a discrete 5-point Likert scale with 0.5-point increments (1=not harmful and 5=very harmful). The annotation interface is shown in Appendix B.2.
@@
As a pilot study, each annotator independently labels a set of 20-30 videos to validate taxonomy coverage, identify potential edge cases, and familiarize themselves with the annotation interface. We monitor disagreements and encourage detailed note‐taking. Annotators then join a group session to discuss flagged issues, collaboratively annotate selected videos while explaining their reasoning, until consensus is reached.
@@
We then flag examples with weak agreement or disagreement on pro-bigorexia detection and examples where severity scores differ by more than two points. The two annotators are invited to a structured discussion to articulate their reasoning, review cases that need consultation, address disagreements, and eventually aim to reach consensus, although consensus is not mandatory. After the discussion, each annotator independently re‐annotates the flagged videos. Further details on this procedure are provided in Appendix  B.4.
@@
While each video may receive multiple label pairs (primary-subcategory), to ensure consistent and reproducible evaluation, we assign a single label pair to each video. For videos with perfect agreement, we randomly choose one of the label pairs. For strong agreement, we adopt the shared primary–subcategory pair; for weak agreement, we choose the mandatory (first) label over the optional (second) one; remaining disagreements are resolved by random tie‐breaking. The final harm severity score is the average of the two annotators’ ratings. This simplification reflects current limitations of LLMs in reliable multi‐label classification Ma et al. (2025).
We assess inter‐rater reliability with Cohen’s $\kappa$  Cohen (1960) for two annotators. Initial agreement is moderate (strong $\kappa=0.43$-$0.69$; weak $\kappa=0.59$-$0.83$; Table 8). After structured consensus discussions and re‐annotation, reliability increases substantially (strong $\kappa=0.58$-$0.81$; weak $\kappa=0.78$-$0.94$), ensuring that our iterative process yields high‐quality, consistent labels. All Cohen’s $\kappa$ values for each batch are shown in Table 8 (Appendix B.5). Severity ratings showed moderate agreement (ICC = 0.413, 69.4% within one point), typical for subjective clinical assessments and highlighting the complexity of harm evaluation in this domain (Figure 6, Appendix B.5).
@@ 4.3 Data Statistics
Primary Category Train Test Total
Relationship to Body 489 98 587
Relationship to Food 310 98 408
Relationship to Exercise 450 98 548
Supplement Abuse 201 98 299
Relationship to Masculinity 66 98 164
Irrelevant 450 98 548
Total 1,966 588 2,554
Subcategory Train Test Total
Muscularity Self‐objectification 170 34 204
Leanness Self‐objectification 59 34 93
Muscle Dissatisfaction 72 34 106
Rigid Food Rules 144 34 178
Unsafe Food 59 34 93
Cheat Meals 60 34 94
Excessive Exercise 129 34 163
Predebting Exercise 28 4 32
Maladaptive Coping 82 34 116
Exercise‐Induced Functional Impairment 43 6 49
Toxic Motivation 102 34 136
Anabolic Steroids 61 34 95
Legal APEDs 56 10 66
Hormone Therapy 77 34 111
Relationship to Masculinity 130 34 164
Irrelevant 200 34 234
Total 1,472 462 1,934
Table 1: BigTok Classification Benchmark statistics showing the final train/test split distribution after data processing.
Our BigTok dataset originates from a raw core set of 2,210 expert-annotated TikTok videos (median duration 20.6 seconds, 59.8% under 30 seconds). To construct robust classification benchmarks, we apply a strict preprocessing pipeline detailed in Section B.4 (Appendix), which includes removing corrupted or non-English files and supplementing the dataset with verified irrelevant videos sampled via popular keywords (Table 7). Table 1 presents the final train/test split statistics resulting from this processing. Specifically, the primary category benchmark (Task 1) comprises 2,554 videos split into 1,966 training and 588 balanced test samples, while the subcategory benchmark (Task 2) utilizes a relevance-filtered subset of 1,934 videos (1,472 training, 462 test). Within the processed primary categories (excluding Irrelevant), Relationship to Body (587 videos) and Relationship to Exercise (548 videos) represent the largest classes. At the subcategory level, Muscularity Self-objectification (204 videos) predominates, reflecting the prevalence of physique-display content on fitness-oriented TikTok. Examples from the final dataset appear in Tables 14 and 15 in Appendix A.1, with additional annotation statistics provided in Appendix B.5. The raw annotation counts are shown in Table 9 in the Appendix.
@@ 5 Classifying Multimodal Pro-Bigorexia Content via VLMs
To develop an automated detection system for pro-bigorexia content, we apply SOTA VLMs to this challenging multimodal classification task. We leverage both proprietary and open-source VLMs through zero-shot prompting, few-shot learning, and finetuning approaches to achieve high performance in identifying subtle pro-bigorexia behaviors across video, audio, and text modalities.
@@ 5.2 Models
Given the multimodal nature of our classification task, involving text from captions, audio transcripts, images, and video content, we require SOTA models designed for multimodal data, particularly video understanding. We evaluate three commercial API‐based VLMs: GPT‑4.1 OpenAI (2025) and Claude-Sonnet-4 Anthropic PBC (2025) (both process images), while Gemini-2.5-Flash Comanici et al. (2025) natively accepts video input. We also evaluate two open-source models: Qwen2.5-VL Qwen Team (2025), and InternVL3 Zhu et al. (2025), which uses Qwen2.5 pre-trained base models and Variable Visual Position, which achieves strong performance on video benchmarks.
Model sizes and versions are listed in Table 10, Appendix D.1. Due to computational constraints, we focus on small and medium-sized models. For open-source models, we set temperature = 0.1 to encourage deterministic outputs. For commercial API models, we utilize the provider defaults (temperature = 1.0). To ensure the validity of this choice, we conducted a robustness check comparing $T=1.0$ vs $T=0.1$ for all commercial models (see Appendix E.1). We observe negligible performance differences, confirming that our results are robust to hyperparameter variations. Implementation details and hardware specifications are provided in Appendix D.1. Details about the train and test datasets are in Appendix C.